Ms. Gebel is the owner of Pink Flower, a local boutique. She stocks two major product lines, a line of jeans and a line of dresses. She wants to understand how the line have performed in the past in terms of GMROI and how future performance can be improved. She has the following information:

Jeans ($)

Dresses ($)

Net sales

200,000

100,000

COGS

100,000

50,000

EOM inventory Jan

15,000

30,000

EOM inventory Feb

20,000

25,000

EOM inventory Mar

15,000

25,000

  1. Using the information in the table above, please calculate the following. Don’t forget to show your work.
    1. Gross margin percentage for both lines
    2. Average inventory at cost for both lines
    3. Sales-to-stock ratio for both lines
    4. GMROI for both lines
    5. Inventory turnover for both lines

  1. If Ms. Gebel wants to improve the GMROI for the line of dresses, what can she do?
  1. Suppose Ms. Gebel is considering several new vendors for a line of menswear. She gives you the information in the table below and asks you to choose a vendor based on the multiattribute analysis method. Please show your work and answer the question below.
    1. Which vendor should she choose? Why?

Issues

Importance

of Issues

Brand A

Brand B

Brand C

Brand D

Vendor reputation

9

6

7

3

5

Merchandise quality

8

6

5

3

9

Product fashionability

7

6

4

8

8

Service

6

6

7

9

5

Promotional assistance

5

6

7

8

5
Answer #1

a)

Gross margin percentage for Jeans : ((net sales-COGS)/net sales)*100 = ((200000-100000)/200000)*100 = 50%
Gross margin percentage for Dresses : ((net sales-COGS)/net sales)*100 = ((100000-50000)/100000)*100 = 50%

b)

Average inventory at cost for Jeans = average(15000,20000,15000) = 16666.66667
Average inventory at cost for Dresses = average(30000,25000,25000) = 26666.66667

c)

Sales-to-stock ratio for Jeans = 200000/16666.66667=12
Sales-to-stock ratio for Dresses = 100000/26666.66667=3.75

d)

GMROI for jeans = Gros margin % * (sales/average inventory) = 50%*(200000/16666.66667) = 6
GMROI for Dresses= Gross margin % * (sales/average inventory) = 50%*(100000/26666.66667) = 1.875